# Karabo: A versatile SKA Observation Simulation Framework

### v1

@@ Abstract
interferometers , data analysis , interferometric techniques[first]organization=Space, Planetary & Astronomical Sciences & Engineering (SPASE), Indian Institute of Technology Kanpur,addressline=Kalyanpur, city=Kanpur, postcode=208016, state=Uttar Pradesh, country=India
[second]organization=Institute for Data Science, FHNW,addressline=Bahnhofstrasse 6, city=Windisch, postcode=5210, country=Switzerland
[third]organization=Institute for Particle Physics and Astrophysics, ETH Zurich,addressline=Wolfgang-Pauli-Strasse 27, city=Zurich, postcode=8009, country=Switzerland
@@ Acknowledgements
We thank Pascal Hitz for running the PINOCCHIO simulation and providing us with the resulting HDF5 file of the HI flux maps used as a sky model to test and develop the HI IM science use case. This work was supported by SERI as part of the SKACH consortium. This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID sk05. Some of the results in this paper have been derived using the healpy and HEALPix packages.

### v2

@@ Abstract
[…]
[third]organization=Institute for Particle Physics and Astrophysics, ETH Zurich,addressline=Wolfgang-Pauli-Strasse 27, city=Zurich, postcode=8093, country=Switzerland
[fourth]organization=Laboratoire d’Astrophysique, Ecole Polytechnique Federale de Lausanne EPFL, adressline=Observatoire de Sauverny, city=Versoix, postcode=1290, country=Switzerland
@@ Acknowledgements
This work was supported in part by the AstroSignals Sinergia grant No. CRII5_193826 from the Swiss National Science Foundation. This work was supported by SERI as part of the SKACH consortium. We also thank Christoph Vögele in the initial stages of the project. This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID sk05. Some of the results in this paper have been derived using the healpy and HEALPix packages.